What is the range of the function f(x)= x^2-4x-12
 Answer by lwsshak3(11628)
       
 What is the range of the function f(x)= x^2-4x-12 **
 To find the range, interchange x and y, then solve for y.
 x=y^2-4y-12
 y^2-4y=x+12
 completing the square
 (y^2-4y+4)=x+12+4
 (y-2)^2=x+16
 Take sqrt of both sides
 y-2=±√(x+16)
 y=±√(x+16)+2
 radican≥0
 x+16≥0
 x≥-16
 Range of f(x):
 [-16,∞)
 | 
  
 | 
 |  |  |